prove or disprove that the complement of integers z in R is an open set


1
Expert's answer
2022-06-07T15:59:27-0400

For each integer "n" , open interval "(n,n+1)" is open and "\mathbb{R}−\mathbb{Z}=⋃_{n∈\mathbb{Z}}(n,n+1)". That is, "\mathbb{R}−\mathbb{Z}" is a union of countable many open sets so "\mathbb{R}−\mathbb{Z}" is open.